OpenStudy (anonymous):

4x=2+y I need to find the slope.

OpenStudy (mathmate):

Slope intercept form gives the slope as m, where y=mx+b If you would transform the equation into the form y=mx+b, you can read off m as the slope. Example: 4x+2y=7 2y=7-4x y=(-4x+7)/2=-2x+7/2 So comparing wih the standard form y=mx+b, m=-2=slope (This is an example).

OpenStudy (anonymous):

m=4 that's your slope.

OpenStudy (anonymous):

4x=y+2 y=mx+c ---------(1) in the above given equation y=4x-2 ---------(2) by comparing equation 1 and 2 we can write m=4